What mass of zinc is required to galvanize the top of a 3.00 m × 5.50 m sheet of iron to a thickness of 0.100 mm of zinc? If the zinc comes from a solution of Zn (NO3)2 and the current is 25.5 A, how long will it take to galvanize the top of the iron? The density of zinc is 7.140 g/cm3.
A. 231 g
B. 240 g
C. 115 g
D. 225 g
1
Expert's answer
2020-06-11T10:28:41-0400

0.100 mm  = 0.0001 m

3.00 m × 5.50 m x 0.0001 m = 0.00165 m3 - volume of Zn

The density of zinc is 7.140 g/cm3 or 7140 kg/m3

So mZn = 0.00165 x 7140 = 11.781 kg = 11781 g

Zn2+ + 2e- = Zn

M (Zn) = 65.39 - molar mass of Zn

n = 2e or 2 x 96500 coulomb (symbolized C) for each 65.39 g

If we have 11781 g: Q = (2 x 96500 x 11781) / 65.39

Q = 34771876.4 C

Q = It

t (time) = Q / I

t = 34771876.4 / 25.5 = 1363602.9 s

= 22726.7 min

= 378.8 hours